for each function find the indicated values...
f(x)= 2x+5
f(6)
f(1)
how do i do this problem?

Answer by stanbon(75887)

for each function find the indicated values...
f(x)= 2x+5
f(6)
f(1)
----
f(6) = 2*6 + 5 = 12+5 = 17
----
f(1) = 2*1 + 5 = 2 + 5 = 7
===============================
Cheers,
Stan H.
==========